The vertices of triangle are [at1t2,a(t1+t2)],[at2t3,a(t2+t3)],
[at3t1,a(t3+t1)]then find the orthocentre of the triangle
 

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
 let the first point be A , second B and third C
Slope of AB=m1= a(t3-t1)/ at2(t3-t1) = 1/t2
Similarly, slope of BC=m2= 1/t3
m1m2= -1 since t2t3 = -1
Therefore this is a right triangle, orthocentre = B
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)